x^(5/4)+1=33 solve
Answer by Alan3354(69443)

x^(5/4)+1=33 solve
----------------
x^(5/4) = 32
x^(1/4) = 2
x = 16
---------
Or,
x^(5/4) = 32
x^5 = 1048576
x = 16
------------
Or if you don't recognize the powers of 2:
(5/4)log(x) = log(32)
log(x) = 4log(32)/5
log(x) = 1.204119983
x = 16